OpenStudy (anonymous):

A bus travels from Oakridge to Hillsboro, a distance of 320 miles. The bus is traveling at 60 mph. The driver is required to contact the station in Hillsboro when he is between 50 and 100 miles from the station. During what time interval should he contact the station at Hillsboro. a. He should contact the station anytime after 3.7 hours. b. He should contact the station anytime after 3.7 hours but before 4.5 hours. c. He should contact the station anytime between 3 and 5 hours. d. He should contact the station anytime before 4.5 hours.

OpenStudy (anonymous):

Set up 2 equations. s = d/t 60 = 220/t 60 = 270/t Then say that it's between these two time periods.

OpenStudy (anonymous):

Just solve for t.

OpenStudy (anonymous):

4.5

OpenStudy (anonymous):

b or d?

OpenStudy (anonymous):

That's one of them, what's the second one?

OpenStudy (anonymous):

What's 60 = 270/t?

OpenStudy (anonymous):

its 4.5

OpenStudy (anonymous):

Oh I mean 60 = 220/t

OpenStudy (anonymous):

It'll come out as an infinite, but round to the tenth place.

OpenStudy (anonymous):

Did you get the answer?

OpenStudy (anonymous):

3.6

OpenStudy (anonymous):

so 3.5?

OpenStudy (anonymous):

i mean 3.7

OpenStudy (anonymous):

When you round, it should become 3.7. Which of the answers involves an interval of 3.7 and 4.5?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

There you go :)

OpenStudy (anonymous):

thank you :)